Problem 3-49 (LO. 8)

During 2019, Chester (a married taxpayer filing a joint return) had the following transactions involving capital assets:

Gain on the sale of an arrowhead collection (acquired as an investment at different times but all pieces have been held for more than 1 year) $6,000
Loss on the sale of IBM Corporation stock (purchased 11 months ago as an investment) (4,000)
Gain on the sale of a city lot (acquired 5 years ago as an investment) 2,000


a. Indicate the tax treatment for each item.

Gain on the sale of an arrowhead collection Long-term capital gain
Loss on the sale of IBM Corporation stock Short-term capital loss
Gain on the sale of a city lot Long-term capital gain

Overall, Chester has a long-term capital gain of $4000. However, $2000 of this gain is from collectibles, which are taxed at a maximum rate of 28% .

If Chester has taxable income (including the property transactions) of $375,000, what is the income tax regarding these transactions?

If Chester has taxable income (including the property transactions) of $75,000, what is the income tax regarding these transactions?
Answer #1

Part A

Income tax = tax on long term capital gain + tax on sale of collectibles = (4000*15%)+(2000*28%) = $1160

(tax rate 15% is applied on long term capital gain because taxable income is from $78,751 - $488,850 for married filing jointly in 2019)

Part B

Income tax = tax on long term capital gain + tax on sale of collectibles = (4000*0%)+(2000*28%) = $560

(tax rate 0% is applied on long term capital gain because taxable income is from $0 - $78750 for married filing jointly in 2019)